[Brahms-l] Shift report 20040101 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 01 2004 - 00:48:53 EST
------------------------------------------------------------
Date: Thu Jan  1 00:48:53 2004, Shift: 16:00-24:00
Supervisor: Oana Ristea, with: 
Log book#15 pages 20 - 20

Shift Summary:
------------------------------------------------------------
1. Beam until 5 pm and from 7.45 pm until now.

2. Run 9288 - D1, D2 = 1/4 A, D5 = 350 A - 71k events (tr1);
   run 9289 - D1, D2, D3, D4 = 1/4 A, D5 = 350 A - 50k events  (tr1);
   run 9290 - non TPC pedestal run;
   run 9291 - D1, D2, D3, D4 = 1/4 A, D5 = 350 A - 198k events  (tr1);
   run 9292 - D1, D2, D3, D4 = 1/4 A, D5 = 350 A - 195k events  (tr1);
   run 9293 - D1, D2, D3, D4 = 1/4 A, D5 = 350 A - 269k events  (tr1).

3. All detectors are up and running.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 1 00:49:03 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 01 2004 - 00:49:22 EST